If three numbers are chosen without replacement from the integers 1, 2, 3, 4, 5, 6, 7, 8, what is the probability that the median value is 3?
a 1/12 b 1/20 c 3/20 d 5/28 e 15/28

Answer by stanbon(75887)

If three numbers are chosen without replacement from the integers
1, 2, 3, 4, 5, 6, 7, 8, what is the probability that the median value is 3?
-----------------------
Ways to succeed:
134,135,136,137,138,234,235,236,237,238: 10 ways
-----
# of possible outcomes: 8C3 = 56
-----------
P(3 in the middle) = 10/56 = 5/28
========================================
Cheers,
Stan H.
========================================

a 1/12 b 1/20 c 3/20 d 5/28 e 15/28